Given that the ground state energy of the hydrogen atom is –13.6 eV, the ground state
energy of positronium (which is a bound state of an electron and a positron) is
1
Expert's answer
2020-09-28T08:10:28-0400

Solution

According to question

Ground state energy of electron is

-13. 6 eV.

Now for positronium case mass will interchanged by reduced mass

Which is given by as below

"m=\frac{m_e\times m_e}{m_e+m_e}=\frac{m_e}{2}"


Relationship between ground state energy of electron and positronium can be expressed as

"E=\frac{-13.6}{2}=6.8eV"

Therefore Ground state energy of positronium is 6.8eV.